# Scalable Vision Language Model Training via High Quality Data Curation

### v1

@@ Abstract
In this paper, we introduce SAIL-VL (ScAlable Vision Language Model TraIning via High QuaLity Data Curation), an open-source vision language model (VLM) of state-of-the-art (SOTA) performance with 2B parameters. We introduce three key improvements that contribute to SAIL-VL’s leading performance: (1) Scalable high-quality visual understanding data construction: We implement a visual understanding data construction pipeline, which enables hundred-million-scale high-quality recaption data annotation. Equipped with this pipeline, we curate SAIL-Caption, a large-scale caption dataset with large quantity and the highest data quality compared with opensource caption datasets. (2) Scalable Pretraining with High-Quality Visual Understanding Data: We scale SAIL-VL’s pretraining budget up to 131B tokens and show that even a 2B VLM benefits from scaled up training data sizes, exhibiting expected data size scaling laws in visual understanding and instruction following performance. (3) Scalable SFT via quantity and quality scaling: We introduce general guidance for instruction data curation to scale up instruction data continuously, allowing us to construct a large SFT dataset with the highest quality. To further improve SAIL-VL’s performance, we propose quality scaling, a multi-stage training recipe with curriculum learning, to improve model performance scaling curves w.r.t. data sizes from logarithmic to be near-linear.
SAIL-VL obtains the highest average score in 19 commonly used benchmarks in our evaluation and achieves top1 performance among VLMs of comparable sizes on OpenCompass (https://rank.opencompass.org.cn/leaderboard-multimodal). We release our SAIL-VL-2B model at HuggingFace (https://huggingface.co/ BytedanceDouyinContent/SAIL-VL-2B).
@@
footnotetext: * Equal contribution.footnotetext: †Email corresponding
@@ 1 Introduction
Researches in vision large language models (VLMs) Liu et al. (2024a); Li et al. (2024a); Yao et al. (2024); Wang et al. (2024); Gu et al. (2024); Chen et al. (2024d, c) have made significant progress in recent years, facilitating various vision tasks via language interactions. Due to the memory and computational constraints in model deployment, training compact VLMs with robust visual comprehension performance has become a popular research field recently Marafioti et al. (2024); Chen et al. (2023a); Yao et al. (2024); Li et al. (2024c); Gao et al. (2024). However, how to make optimal use of publicly available resources to unlock the potential of compact VLMs remains an unanswered question. We attribute the suboptimal performance of recent lightweight vision language models to two aspects: limited fundamental visual understanding abilities and unsatisfactory visual instruction following performance.
@@
The fundamental visual understanding abilities of VLMs are typically established via large-scale pretraining. VLM pretraining necessitates not only substantial training budgets, but also a sufficient amount of high-quality visual understanding data to take effect. Recently proposed VLMs, such as LLaVA series Liu et al. (2024a); Li et al. (2024a); Chen et al. (2024a), conduct light-weight pretraining with a limited amount of low-quality caption data, and therefore suffer from suboptimal visual understanding abilities which hinder subsequent visual instruction tuning. MiniCPM-V-2.5 Yao et al. (2024) and Qwen2-VL Wang et al. (2024) allocate hundreds of billions of tokens’ computation budgets to the pretraining stage, but the limited visual understanding data quality hinders these models from achieving stronger visual understanding capabilities. More importantly, despite the large amount of resources consumed in VLM pretraining, existing works do not provide reliable conclusions to understand how pretraining budgets and data quality influence VLM performance.
@@
During supervised fine-tuning (SFT) stage, VLM’s visual understanding capabilities are generalized to instruction following tasks. However, how to make optimal use of high-quality visual instruction tuning datasets remains unexplored. Recent works primarily focus on adjusting the data distribution across various domains and formats Li et al. (2024a); Chen et al. (2024d, c); Yao et al. (2024). Infinity-MM Gu et al. (2024) proposes multi-stage SFT to improve model performance, and obtain promising performance scaling results. Nevertheless, there still lacks widely acknowledged guidelines to determine the distribution of SFT datasets and training stages.
@@
To address the above issues, we propose SAIL-VL, an opensource vision language model with state-of-the-art (SOTA) visual understanding and instruction following performance. SAIL-VL is trained through five stages: pretrain-alignment, pretrain-advance, SFT-knowledge, SFT-instruction and SFT-preference. In the first two pretraining stages, we establish SAIL-VL’s fundamental visual understanding abilities with large-scale pretraining. To explore how pretraining computation budgets and data quality influence VLM’s visual understanding abilities, we scale up VLM pretraining to 131B tokens with SAIL-Caption, our synthesized visual understanding dataset, which is of large magnitude and the highest quality compared to other publicly available caption datasets. During the following three SFT stages, we train SAIL-VL through three separate stages with different training objectives. Moreover, we elaborate guidelines to select SFT datasets with balanced distribution and the guidelines for the design of VLM SFT stages to further improve model performance. Trained through five stages, our SAIL-VL-2B achieves top-1 performance on OpenCompass leaderboard compared to VLMs of similar sizes, and obtains the highest average score in 19 commonly used benchmarks in our evaluation.
@@
(1) We implement a data construction pipeline for scalable high-quality visual understanding data construction. Equipped with this pipeline, we construct SAIL-Caption dataset, which is of large magnitude and the highest quality compared with opensource caption datasets.
(2) We scale up SAIL-VL’s pretraining data size to 131B tokens, and report logarithmic model performance scaling laws w.r.t. training data size and training budgets. To the best of our knowledge, this is the first time that data size scaling laws for VLM pretraining are proposed and discussed.
(3) We elaborate the guidelines for the design of VLM SFT strategies and the distribution of visual instruction tuning datasets, which enable SAIL-VL-2B to achieve top1 performance on OpenCompass leaderboard and obtain the highest average score in 19 commonly used VLM benchmarks compared with previous SOTA models of comparable sizes.
@@ 2 Model Training Pipeline
In this section, we introduce SAIL-VL’s training strategy shown in Fig 1. Starting from InternViT-300M Chen et al. (2023b) and Qwen-2.5-1.5B Team (2024b), SAIL-VL is pretrained for visual understanding in the first two pretraining stages and then adapted to visual instruction tasks in instruction tuning stages.
Figure 1: SAIL-VL’s overall model training pipeline.
2.1 Pretrain
During the pretraining stages, we gradually open model parameters for training to develop SAIL-VL’s fundamental visual understanding abilities from large-scale pretraining data.
Pretrain-Alignment.
We start from a randomly initialized multi-layer perceptron (MLP) module which bridges the representation spaces of visual features and language models. We train the vision-to-language projector with approximately 16B tokens in this stage, equipping SAIL-VL with basic visual understanding abilities. Detail caption data and OCR data are used to train the model in this stage.
Pretrain-Advance.
After the alignment pretrain stage, we unlock the visual encoder of SAIL-VL and train the model with approximately 131B tokens. We include more OCR data and a larger proportion of SAIL-Caption for model training. We do not use the entire SAIL-Caption dataset but a subset of it, selected to maintain an even distribution across different data sources, thereby ensuring diversity in data distribution.
2.2 Instruction Tuning Stage (SFT)
We train all parameters of SAIL-VL through three SFT stages in a curriculum learning fashion with progressively higher-quality training data.
SFT-Knowledge.
In this stage, SAIL-VL learns basic instruction-following abilities and ingests world knowledge from the training data. We use the data from Infinity-MM Stage2 Gu et al. (2024) for SFT-Knowledge stage training, which is of the largest quantity and medium quality among training datasets for the three SFT stages.
SFT-Instruction.
We further optimize the SAIL-VL model to obtain enhanced visual instruction following capabilities with a balanced distribution across downstream domains. We use a 12M customized visual instruction tuning dataset, including more open-ended question-answering, complex OCR VQA, and other visual instruction following data with more diverse distribution, prioritizing data quality over sheer volume.
SFT-Preference.
During the last stage, we train SAIL-VL on scarce yet high-quality data with open-ended answers aligned with human preference, including LLaVA Li et al. (2024a) SFT data, Molmo Caption Deitke et al. (2024), and Infinity-MM Stage4 Gu et al. (2024). Preference learning data for this stage is more complex and harder to collect, and is therefore used in the last SFT stage for training effectiveness.
3 Towards Scalable VLM Training
Dataset Language # Sample Avg. Len. Uni. 1-gram Uni. 2-gram Uni. 3-gram Uni. Noun Uni. Verb Uni. Adj.
@@ 3 Towards Scalable VLM Training
DataComp-LLaVA-Caption (2024b) EN 940,891,257 48.37 1.661 8.400 18.23 0.4728 0.0333 0.2801
@@
Table 1: Statistics of SAIL-Caption and other opensource caption datasets. We employ NLTK Bird (2006) and Jieba Sun to perform text segment and part-of-speech tagging for English and Chinese captions, respectively. “Avg. Len.” stands for “average length” and “Uni.” denotes “unique”.
We elaborate the design of our scalable model training pipeline: We first establish a scalable high-quality data construction pipeline, and then validate the data scaling laws observed in pretraining with this data. Additionally, we introduce the scaling law for instruction tuning, a data curation pipeline for quantity scaling, and a multi-stage training pipeline for quality scaling.
@@ 3.1 Scalable High-Quality Visual Understanding Data Construction
We illustrate our scalable high-quality detail caption data construction pipeline in Figure 1, which can be described as the following four steps.
@@
We collect source data from a wide range of publicly available datasets to ensure our visual understanding datasets possess sufficiently diverse distribution. Our source datasets include LAION-COCO Schuhmann et al. (2022), TextCaps Sidorov et al. (2020), SA1B Kirillov et al. (2023), and several other large-scale datasets.
@@
We curate a small amount of reference data to train an efficient VLM to annotate high-quality visual understanding data efficiently at scale. We first select a subset of our collected source images with a balanced distribution across different source datasets, and then task powerful VLM APIs to annotate detail captions. It is worth noticing that we follow previous works Yu et al. (2024); Hong et al. (2024) to provide alt-text for reference data generation if possible, providing supplementary world knowledge for enhanced reference data quality.
@@
Equipped with the high-quality reference data, we train a compact VLM for scalable caption data generation. We train SAIL-Captioner on the reference dataset, with alt-text optionally included in the prompt, enabling SAIL-Captioner to perform both captioning and recaptioning.
@@
In the final stage, we use SAIL-Captioner to annotate large-scale visual understanding datasets for model pretraining. We implement a data construction pipeline, which supports multi-task, multi-node, and multi-processing asynchronous annotation, to enable flexible data annotation resource allocation. LMDeploy Contributors (2023) is incorporated in our annotation system to further speed up the generation process.

### v3

@@ Abstract
In this paper, we introduce SAIL-VL (ScAlable Vision Language Model TraIning via High QuaLity Data Curation), an open-source vision language model (VLM) series achieving state-of-the-art (SOTA) performance in 2B and 8B parameters. The following three key improvements contribute to SAIL-VL’s leading performance: (1) Scalable high-quality visual understanding data construction: We implement a data construction pipeline to enable hundred-million-scale high-quality recaption data annotation. The resulted dataset SAIL-Caption is validated to be of the highest data quality compared with opensource datasets. (2) Scalable Pretraining with High-Quality Visual Understanding Data: We scale SAIL-VL’s pretraining budget up to 655B tokens and show that even a 2B VLM benefits from scaled up training data sizes, exhibiting logarithmic data size scaling laws in benchmark performance. (3) Scalable SFT via data quantity and complexity scaling: We curate a high-quality SFT dataset collection with leading data quantity scaling effectiveness and demonstrate that training with progressively higher-complexity data surpasses baseline one-stage training by a large margin.
SAIL-VL series models achieve the highest average score in 18 widely used VLM benchmarks in our evaluation, with the 2B model takes the top position over VLMs of comparable sizes on OpenCompass 2024 (https://rank.opencompass.org.cn/leaderboard-multimodal), demonstrating robust visual comprehension abilities. SAIL-VL series models are released at HuggingFace (https:// huggingface.co/BytedanceDouyinContent).
@@
* Equal contribution.†Email corresponding
Figure 1: SAIL-VL’s overall data construction and model training pipeline, as well as data size scaling laws observed in our large-scale VLM training experiments.
@@ 1 Introduction
Researches in large vision language models (VLMs) Liu et al. (2024a); Li et al. (2024a); Yao et al. (2024); Wang et al. (2024b); Gu et al. (2024); Chen et al. (2024d, c) have made significant progress in recent years, facilitating various vision tasks via language interactions. Due to the memory and computational constraints in model deployment, training compact VLMs with robust visual comprehension performance has become a popular research field recently Marafioti et al. (2024); Chen et al. (2023a); Yao et al. (2024); Li et al. (2024c); Gao et al. (2024). However, how to make optimal use of publicly available resources to unlock the potential of compact VLMs remains an unanswered question. We attribute the suboptimal performance of recent lightweight vision language models to their limited fundamental visual understanding abilities and unsatisfactory instruction following performance.
@@
The fundamental visual understanding abilities of VLMs are typically established via large-scale pretraining, which necessitates not only substantial training budgets, but also a sufficient amount of high-quality visual understanding data to take effect. Recently proposed VLMs, such as LLaVA series Liu et al. (2024a); Li et al. (2024a); Chen et al. (2024a), conduct light-weight pretraining with a limited amount of low-quality caption data, and therefore suffer from suboptimal visual understanding abilities which hinder subsequent visual instruction tuning. MiniCPM-V-2.5 Yao et al. (2024) and Qwen2-VL Wang et al. (2024b) allocate hundreds of billions of tokens’ computation budgets to the pretraining stage, but the limited visual understanding data quality undermines their visual understanding performance. More importantly, despite the large amount of resources consumed in pretraining, existing works do not provide reliable conclusions to understand how pretraining budgets and data quality influence VLM performance.
@@
During the supervised fine-tuning (SFT) stage, VLM’s visual understanding capabilities are generalized to instruction following tasks. However, how to make optimal use of high-quality visual instruction tuning datasets remains unexplored. To obtain SFT data collections with higher quality, recent works focus on adjusting the data distribution across various domains and formats Li et al. (2024a); Chen et al. (2024d, c); Yao et al. (2024). Infinity-MM Gu et al. (2024) further explores enhancing the data efficiency of visual instruction tuning datasets with a multi-stage SFT strategy, obtaining promising performance scaling results. Despite the promising results of these works, there still lacks widely acknowledged methodologies to determine the distribution of SFT dataset collections or allocation of SFT stages.
@@
To address the above issues, we propose SAIL-VL, an opensource vision language model series in 2B and 8B parameters with state-of-the-art (SOTA) performance. SAIL-VL is trained through several pretraining and SFT stages. We first establish SAIL-VL’s basic visual understanding abilities via large-scale pretraining. To explore how pretraining computation budgets and data quality influence VLM performance, we scale up VLM pretraining to 655B tokens with SAIL-Caption, our synthesized large-scale detail caption dataset with top data quality compared to opensource alternatives. During the following SFT stages, we train SAIL-VL on our customized SFT data collection which outperforms opensource datasets markedly in data quality. SAIL-VL is trained in a curriculum learning paradigm of three stages, leading to improved data efficiency and model performance. The resulting SAIL-VL-2B and 8B models achieve new SOTA performance in 18 widely used VLM benchmarks.
@@
(1) We implement a data construction pipeline for scalable high-quality visual understanding data construction, equipped with which we construct SAIL-Caption, which is of large quantity and the highest quality compared with opensource datasets.
(2) We scale up SAIL-VL’s pretraining data size to 655B tokens, and report logarithmic model performance scaling laws w.r.t. training data sizes. To the best of our knowledge, this is the first time that data size scaling laws for VLM pretraining are proposed and discussed.
(3) We elaborate on the methodologies for high-quality SFT data curation, and demonstrate the effectiveness of the curriculum SFT strategy. Our SAIL-VL series models achieve top-ranked performance in our evaluation on 18 opensource VLM benchmarks.
@@ 2 Model Training Pipeline
In this section, we introduce SAIL-VL’s training strategy as shown in Fig 1. Starting from InternViT Chen et al. (2023b) and Qwen-2.5 Team (2024b) series models, SAIL-VL is pretrained for visual understanding and adapted to instruction following tasks in a total of five training stages.
Dataset Language # Sample Avg. Len. Quality Uni. 2-gram Uni. 3-gram Uni. Noun Uni. Verb Uni. Adj.
@@
DataComp-LLaVA-Caption (2024b) EN 940,891,257 48.37 70.0 8.400 18.23 0.4728 0.0333 0.2801
@@
BLIP3-KALE (2024) CN 235,125,090 66.16 73.2 21.08 42.83 0.8686 0.0930 0.7012
[…]
Table 1: Statistics of SAIL-Caption and other opensource datasets. “Quality” refers to quality scores evaluated by human annotators. We employ NLTK Bird (2006) and Jieba Sun to perform text segmentation and part-of-speech tagging for English and Chinese captions, respectively. “Avg. Len.” stands for “average length” and “Uni.” denotes “unique” items per sample. Statistics of SAIL-Caption subsets are marked with .
2.1 Pretrain
During pretraining, we gradually open model parameters for larger-scale pretraining to develop SAIL-VL’s visual understanding abilities. We start from a randomly initialized multi-layer perceptron (MLP) module as the vision-to-language projector, and train it with approximately 131B tokens of detail caption and OCR data in the Pretrain-Alignment stage. After warming up, we unlock the visual encoder of SAIL-VL for larger model capacity during the following Pretrain-Advance stage, and train the model through approximately 524B tokens. Note that we do not use the entire SAIL-Caption dataset but a subset with an even distribution instead to ensure the diversity in data distribution. For OCR data, we use several high-quality OCR datasets repeatedly instead of incorporating diverse but relatively low-quality data. The advantage of using repeated-yet-high-quality data is shown in Section 5.1. For SAIL-VL-8B, we allocate 20B- and 32B-token training budgets in the two pretraining stages for efficiency.
2.2 SFT
We train all parameters of SAIL-VL in a curriculum learning fashion with progressively higher-complexity training data in SFT stages. In the first SFT-Knowledge stage, SAIL-VL learns basic instruction-following abilities and ingests world knowledge from Infinity-MM Stage2 Gu et al. (2024) data. During the subsequent SFT-Instruction stage, we further optimize SAIL-VL towards enhanced visual instruction following capabilities with our customized 12M-sample high-quality visual instruction tuning dataset. For the final SFT-Preference stage, we train SAIL-VL on a small amount of complex visual instruction tuning data, including LLaVA Li et al. (2024a) SFT, Molmo Caption Deitke et al. (2024), and Infinity-MM Stage4 Gu et al. (2024) data, enabling SAIL-VL to tackle a wider range of complex instruction following tasks. We refer to Section 5.2 for detailed data distribution of the three stages.
3 Towards Scalable VLM Training
In this section, we introduce our scalable high-quality data construction pipeline and elaborate on the model performance scaling laws observed in both pretraining and SFT stages.
@@ 3.1 Scalable High-Quality Visual Understanding Data Construction
Our scalable data construction pipeline is shown in Figure 1, consisting of the following four steps.
@@
We collect source data from a wide range of public image datasets to ensure data distribution diversity. Our source datasets include LAION-COCO Schuhmann et al. (2022), TextCaps Sidorov et al. (2020), SA1B Kirillov et al. (2023), and several other large-scale datasets.
@@
We curate a small amount of reference data to train a compact VLM for efficient data annotation at scale. We first select a subset of source images with a balanced distribution, and then task GPT4-O-20240513 OpenAI (2024) deployed by Azure to annotate detail captions. Following previous works Yu et al. (2024a); Hong et al. (2024), alt-texts are provided if available for supplementary world knowledge and enhanced reference data quality.
@@
Equipped with the high-quality reference data, we train an InternVL2-8B Team (2024a) model on the reference data to generate high-quality data at scale, which is called SAIL-Captioner. Similarly, alt-texts are optionally included in the caption generation prompt, enabling SAIL-Captioner to perform both captioning and recaptioning tasks.
@@
In the final stage, we deploy SAIL-Captioner with LMDeploy Contributors (2023) for large-scale detail caption data construction. We implement a multi-task, multi-node, and multi-processing asynchronous annotation pipeline, enabling flexible computation resource allocation.
Figure 2: Scaling curves of SAIL-VL-2B’s performance dynamics in the pretrain-alignment (PT-Ali) stage. We show model performance on all understanding benchmarks, caption tasks and OCR tasks, respectively. “BMK Score” stands for average benchmark scores.
Figure 3: Scaling curves of SAIL-VL-2B’s performance dynamics in the pretrain-advance (PT-Adv) stage. We show pretrained and SFT model performance on understanding benchmarks and OS (opensource) VLM benchmarks, respectively. “BMK Score” stands for average benchmark scores.
@@
Equipped with the aforementioned data construction pipeline, we construct SAIL-Caption, a detail caption dataset with 300M image samples from various sources. To validate the data quality of SAIL-Caption, we randomly sample 10,000 cases from SAIL-Caption and other opensource caption datasets for comparison, and the statistics are shown in Table 1. Results show that SAIL-Caption is not only of large quantity, but also demonstrates leading richness of visual elements, for example, unique n-grams, nouns, verbs, and adjectives in caption texts. These statistics indicate that SAIL-Caption encompasses more visual elements and exhibits greater linguistic diversity in caption texts. Moreover, SAIL-Caption receives higher quality scores from human annotators, surpassing existing opensource datasets by a large margin. We refer to Appendix D for detailed caption quality evaluation procedure and SAIL-Caption showcases.
In this part, we introduce the data size scaling laws observed in SAIL-VL-2B large-scale pretraining. For model checkpoints obtained at different pretraining steps, we conduct lightweight annealing training with 2M identically distributed data for improved convergence and evaluation stability.
SAIL-VL-2B is trained through 131B and 524B tokens during the two pretraining stages, respectively, during which we investigate model performance dynamics. To evaluate the visual understanding performance of SAIL-VL, we establish an evaluation suite which covers fundamental visual understanding tasks such as detail caption generation Dong et al. (2024) and OCR detection Biten et al. (2022); Wang et al. (2020); Gupta et al. (2016); Kim et al. (2022). Details can be found in Appendix E.
As shown in Figure 2, SAIL-VL’s visual understanding performance in each domain improves steadily in the pretrain-alignment stage. As the training data size scales up exponentially, the model performance exhibits a linear growth trend. We also show the understanding performance dynamics in the pretrain-advance stage. SAIL-VL’s understanding benchmark scores improve markedly in this stage, which we attribute to the large capacity of the vision encoder optimized for visual understanding. In Figure 3 (a), a similar linear performance scaling curve is observed, unveiling a promising prospect to scale up VLM pretraining data sizes for improved model performance.
To further investigate the effectiveness of SAIL-VL’s large-scale pretraining, we conduct SFT with different data collections for pretrain-advance model checkpoints trained with different data sizes.
As shown in Figure 3 (b)(c), the overall performance dynamics of SFT models can be plotted as a near-linear curve on an exponential horizontal axis, exhibiting smooth data size scaling laws on opensource VLM benchmarks. We conduct experiments with both our SFT-Instruction data and opensource LLaVA-Next SFT data. Despite the different data composition and final benchmark scores, similar scaling curves can be observed in both experiment sets. We further investigate pretrained and SFT model performance correlation in Appendix G.
Figure 4: Scaling curves of model performance trained on our SAIL-Instruct dataset, LLaVA-OneVision Li et al. (2024a) single image SFT data, and datasets from Infinity-MM Gu et al. (2024). Model performance is shown as an average score across 18 benchmarks.
3.3 Scaling up Visual Instruction Tuning
Despite the abundance of publicly available visual instruction tuning data, high-quality training data is still scarce. We first introduce guidelines for our high-quality SFT data curation in Section 3.3.1, and demonstrate model performance scaling laws of the curriculum SFT strategy in Section 3.3.2.
3.3.1 High-Quality SFT Data Curation for Data Quantity Scaling
In this part, we elaborate on the methodologies for visual instruction tuning data curation and demonstrate their effectiveness in SAIL-VL training.
Figure 5: Model performance dynamics of the quality scaling and all-in-one (AIO) training strategy. “AIO learning” incorporates all three-stage SFT data into a single training loop. Model performance is shown as an average score across 18 benchmarks.
High-quality visual instruction tuning dataset curation.
To judge the quality of different SFT data collections efficiently, we start with the Quick Quality Evaluation strategy. This strategy assesses the quality of a given SFT data collection by training with its 2M-sample subset. The resulting model performance reflects the training data quality, enabling efficient data quality evaluation and comparison. In this strategy, we assume that models trained on different datasets maintain a consistent performance ranking across varying training data sizes. This assumption is validated by experiment results shown in Figure 4.
SFT Stage Size $\downarrow$ Diff. $\uparrow$ Comp. $\uparrow$ Rel. $\uparrow$
SFT-Knowledge 21M 1.90 2.44 3.94
SFT-Instruction 12M 2.15 2.62 4.45
SFT-Preference 3.5M 2.20 2.74 4.55
Table 2: Sizes and quality evaluation results of the three-stage SFT data. “Diff.”, “Comp.”, and “Rel.” stand for task difficulty, data complexity, and imaget-text relevance, respectively.
Pretrain-Alignment Pretrain-Advance SFT-Knowledge SFT-Instruction SFT-Preference
2B 8B 2B 8B 2B 8B 2B 8B 2B 8B
Vision Resolution $448\times\{\{1\times 1\},...,\{2\times 4\}\}$ $448\times\{\{1\times 1\},...,\{2\times 4\}\}$ $448\times\{\{1\times 1\},...,\{2\times 5\}\}$ $448\times\{\{1\times 1\},...,\{2\times 5\}\}$ $448\times\{\{1\times 1\},...,\{2\times 5\}\}$
# Max Visual Token 2048 2048 2560 2560 2560
Data Data Composition SAIL-Caption & OCR SAIL-Caption & OCR Curated VQA Data Curated VQA Data Curated VQA Data
Dataset Size 64M 10M 256M 16M 21M 12M 3.5M
Training Trainable Module Projector Vision & Projector Full Model Full Model Full Model
Trainable Parameter 8.65M 27.52M 313M 332M 1.85B 7.95B 1.85B 7.95B 1.85B 7.95B
Batch Size 1920 512 2048 2048 512 512 512
Learning Rate $1\times 10^{-4}$ $1\times 10^{-3}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $2\times 10^{-5}$ $1\times 10^{-5}$ $2\times 10^{-5}$ $1\times 10^{-5}$ $2\times 10^{-5}$ $1\times 10^{-5}$
Table 3: Details of the training pipeline of SAIL-VL-2B and SAIL-VL-8B.
We then propose the Composition Evaluation strategy to judge the quality of existing SFT data components. In composition evaluation, we start with existing SFT data collections, for example, LLaVA-OneVision Li et al. (2024a), Cauldron laurenccon2024matters. We then categorize the datasets based on their format and distribution, resulting in a series of data groups, including closed-form VQA, open-ended VQA, document VQA, math&reasoning, and pure text QA data  [^1] Closed-form and open-ended VQA data refer to natural image VQA data requiring specific choice answers and open-ended responses, respectively. To optimize the proportion of different data components, we halve each data group and judge the quality of the resulting data collection with our quick quality evaluation method. Once the model performance improves, the downward adjustment of the data proportion is retained.
For incoming datasets to be incorporated into the SFT data, we conduct Incremental Evaluation. Each new dataset is included in the SFT data collection, with the resulting data quality evaluated via lightweight model training. Datasets improving the model performance are regarded as beneficial for the data quality, and are therefore incorporated into our data collection. We also incorporate datasets which maintain the model performance, as they help expand the data scale for improved results.
Data Quantity Scaling.
We curate our SAIL-Instruction data collection (used in SFT-Instruction stage) with the methodologies described above. To validate its advantage in data quality, we train the SAIL-VL model with our SAIL-Instruction data and other opensource SFT data collections at varying data scales. As shown in Figure 4, the performance of SAIL-VL scales up stably as the model training proceeds, depicting a logarithmic performance scaling curve. Compared with other opensource SFT data collections, our SAIL-Instruction data achieves the highest model performance at every data point. It is also worth noticing that the performance ranking of models trained with different datasets remains consistent across the training process. This observation validates our quick quality evaluation method introduced above.
3.3.2 Multi-stage Instruction Tuning for Data Complexity Scaling
In this part, we introduce data complexity scaling, a curriculum learning strategy for VLM SFT for enhanced model performance.
Curriculum SFT with progressively improving data quality.
As elaborated in Section 2.2, we train SAIL-VL through three SFT stages, and the data collections used in later stages differ from previous ones in the following aspects: (1) Datasets are harder to collect and therefore of smaller quantity. (2) Training tasks become increasingly challenging, and the questions in the training data are more difficult to answer. (3) Data complexity progressively increases, requiring more fine-grained understanding of the visual elements and in-depth reasoning. We validate our design by quantifying the data distribution variance across the three stages via human evaluation. As shown in Table 6, the task difficulty, data complexity, and image-text relevance increase monotonously across the three stages. SFT data in the later stages is of higher overall quality, but is also more challenging for the model to learn from, which coincides with our curriculum SFT design. We refer to Appendix F for the detailed definition of these data quality dimensions and the full instruction for human evaluation.
Data Complexity Scaling.
To demonstrate the effectiveness of our curriculum SFT strategy with progressively higher-complexity data, we show model performance dynamics derived from the three SFT stages in comparison with an all-in-one (AIO) training strategy in Figure 5. The model trained with our curriculum SFT strategy exhibits a near-linear performance scaling curve across training stages, outperforming the logarithmic scaling curve of AIO training baseline. This result validates the marked effectiveness of the curriculum SFT strategy. Training with small and high-complexity SFT data in later stages yields more promising performance scaling curves.
Benchmark SAIL-VL Qwen2-VL InternVL2.5-MPO DeepSeekVL-2 SAIL-VL Qwen2-VL InternVL2.5-MPO DeepSeekVL-2
2B Model 8B Model
Overall Performance
Opensource Average 69.1 64.4 67.7 67.0 74.5 73.0 74.3 72.7
   General VQA 60.4 58.3 63.1 59.4 68.3 68.5 71.2 66.8
   OCR VQA 75.9 72.5 71.1 74.4 79.8 79.6 76.3 79.0
   Math&Knowledge 79.0 59.0 75.3 71.3 83.3 71.0 83.2 79.0
   Hallucination 66.2 62.9 64.5 63.6 68.7 67.5 69.7 65.3
General VQA
MMStar (2024b) 55.1 46.3 54.3 49.9 64.2 58.3 65.3 57.7
$\text{MMBench}_{\text{DEV}}$ (2024c) 72.4 68.8 72.5 68.3 79.5 79.5 83.3 78.1
$\text{MMMU}_{\text{VAL}}$ (2024) 40.1 39.9 41.2 39.6 48.2 50.9 52.8 47.6
MME (2023) 1969 1923 2123 1910 2244 2321 2321 2149
$\text{SEEDBench}_{\text{IMG}}$ (2023a) 74.7 72.0 73.2 72.5 75.5 75.3 76.9 76.8
RealWorldQA (2024) 63.8 60.9 60.7 64.8 71.9 69.7 70.2 70.2
MMVet (2024b) 46.1 51.2 64.0 52.8 58.3 62.6 66.8 60.3
OCR VQA